Given the following five-number summary, find the IQR.
2.9, 5.7, 10.0, 13.2, 21.1
The answer is 7.5, but how do I get that?

Answer by rothauserc(4718)

The IQR is the difference between the third quartile and the first quartile
:
minimum is 2.9
:
median is the number which half of numbers are below and half of the numbers are above (2.9, 5.7, 10.0, 13.2, 21.1) = 10.0
:
first quartile is the median of the data below the overall median = (2.9 + 5.7) / 2 = 4.3
:
third quartile is the median of the data above the overall median = (13.2 + 21.1) / 2 = 17.15
:
****************************
IQR = 17.15 - 4.3 = 12.85
****************************
: